# Splitting in a complete local ring and decomposition its group of units

### v4

@@ Abstract.
Let $(R,M,k)$ be a complete local ring (not necessarily Noetherian). Then we reprove by a new method that the natural surjective ring map $R\rightarrow k$ admits a splitting if and only if $\operatorname{char}(R)=\operatorname{char}(k)$. In our proof there is no need for the existence of the coefficient field for equi-characteristic complete local rings, whose existence is the hardest part of the known proof.
However, the main result of this article is that in the unequal characteristic case $\operatorname{char}(R)\neq\operatorname{char}(k)$, we prove that the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
As an application of the above theorem, we show that for any complete local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@
is always split. Next, we show with an example that the above exact sequence does not split for many incomplete local rings.
@@ 1. Introduction
In mathematics, it is common for a particular case of a hidden fact to appear first of all. Then, as the obstacles and coverings are gradually removed, the general form of that phenomenon becomes fully revealed. The same happened in this research work. More precisely, by elementary methods we first showed that for any “finite” local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@
is split. Next, this observation naturally led us to the question of whether the same splitting holds for arbitrary (infinite) local rings.
@@
The first part of this theorem is well known. In fact, it easily follows from the existence of the coefficient field (see e.g. [8, Theorem 28.3(ii)]). Although the existence of the coefficient fields for equi-characteristic complete local rings is a stronger result, proving their existence is the most difficult part of the proof. However, the second part of the above theorem is a new result.
In our proof there is no need for the existence of the coefficient field. The characteristic of the residue field $k=R/M$ plays an important role in our proof and hence it is necessary to examine it carefully. We know that in any local ring $(R,M,k)$, if $\operatorname{char}(R)=\operatorname{char}(k)$ then there are two cases: $\operatorname{char}(k)=0$ or $\operatorname{char}(k)=p$ for some prime number $p$. However, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the characteristic of $k$ is always a prime number $p$.
To prove the above theorem, in the equi-characteristic case, one of the main ingredients that applied is the following important result:
@@
Splitting in a finite local ring is presented in Section 2. But splittings in a general complete local ring is presented in §3.
@@ 3. The general case
It is also clear that if a local ring admits a coefficient field then it is equi-characteristic. As mentioned earlier, the converse is also true for complete local rings: Every equi-characteristic complete local ring has a coefficient field. For its proof see e.g. [8, Theorem 28.3(ii)].

### v5

@@ Abstract.
Let $(R,M,k)$ be a complete local ring (not necessarily Noetherian). As the first main result of this article, we prove that in the unequal characteristic case $\operatorname{char}(R)\neq\operatorname{char}(k)$, the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
Next, we reprove by a new method that in the equi-characteristic case $\operatorname{char}(R)=\operatorname{char}(k)$, the natural surjective ring map $R\rightarrow k$ admits a splitting. In our proof there is no need for the existence of the coefficient fields for equi-characteristic complete local rings, whose existence is the most difficult part of the known proof.
As an application, we show that for any complete local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@
is always split. In particular, we have an isomorphism of Abelian groups $R^{\ast}\simeq(1+M)\times k^{\ast}$. We also show with an example that the above exact sequence does not split for many incomplete local rings.
@@ 1. Introduction
In this article, we first show that for any “finite” local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@
is split. Next, this observation naturally led us to the question of whether the same splitting holds for any “infinite” local ring.
@@
The first part of this theorem is well known. In fact, it easily follows from the existence of coefficient fields (see e.g. [8, Theorem 28.3(ii)]). The existence of coefficient fields for equi-characteristic complete local rings is the most difficult part of this known proof. However, our proof does not require the existence of the coefficient fields. The second part of the above theorem is a new result.
The characteristic of the residue field $k=R/M$ plays an important role in our proof and hence it is necessary to examine it carefully. We know that in any local ring $(R,M,k)$, if $\operatorname{char}(R)=\operatorname{char}(k)$ then there are two cases: $\operatorname{char}(k)=0$ or $\operatorname{char}(k)=p$ for some prime number $p$. However, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the characteristic of $k$ is always a prime number $p$.
To prove the above theorem in the equi-characteristic case, one of the main ingredients that applied is the following important result:
@@
Splitting in a finite local ring is presented in Section 2. But splitting in a general complete local ring is presented in §3.
@@ 3. The general case
It is also clear that if a local ring has a coefficient field then it is equi-characteristic. As mentioned earlier, the converse is also true for complete local rings:
Corollary 3.26.
For a complete local ring $(R,M,k)$ the following statements are equivalent:
$\mathbf{(i)}$ $R$ is equi-characteristic, i.e., $\operatorname{char}(R)=\operatorname{char}(k)$.
$\mathbf{(ii)}$ The natural ring map $R\rightarrow k$ admits a splitting.
$\mathbf{(iii)}$ $R$ has a coefficient field.
Proof.
[…]